# Extracting memorized pieces of (copyrighted) books from open-weight language models

### v3

@@ 1 Introduction
In the dozens of copyright lawsuits over training large language models (LLMs), both sides tend to present opposing, simplified interpretations of the underlying technology. Plaintiffs—such as book authors—say that LLMs are giant copy machines that store (infringing) copies of their specific works and recombine them in outputs [71]. Defendants—typically tech companies—claim that LLMs merely contain non-expressive “statistical correlations” [27], and therefore do not copy any particular creative work. We show that the situation is more complicated than either side suggests.
@@
During training, an LLM processes pieces of text in its training data and learns probabilistic relationships across the dataset as a whole. This informal intuition aligns with the “statistical correlations” that defendants describe as encoded in an LLM’s weights. However, sometimes, for some training data, these learned patterns are not so high-level or abstract; rather, they are highly specific [28]. For reasons that are not fully understood, certain training data are not so transformed by training: they are “memorized” by the model—encoded in some form inside the LLM’s weights [18, 76, 28, 15]. Such memorized content can sometimes be “extracted” later; they can be reproduced in an LLM’s outputs at generation time [16, 60, 93].
@@
Models do not memorize all of their training data, but they do memorize some of it. An LLM may therefore memorize parts of some plaintiffs’ copyrighted works, but not others’. When such memorization occurs, it involves copying that may have significant consequences for ongoing copyright litigation. Extraction generates a “copy” of training data, but it also demonstrates the existence of a “copy” of that training data is memorized inside the model itself. The model being a “copy” (in a technical sense that copyright cares about) has important implications. Notably, a model—not just extracted training data—could be deemed an infringing “copy” of the training data it has memorized [28, 76]. Copyright law offers the destruction of infringing materials as a remedy. So, just as courts have ordered the destruction of bootleg DVDs, they could order the destruction of infringing models [76, 111, 132].
@@
When seeking such remedies, plaintiffs often cite technical papers that demonstrate memorization by extracting training data from open-weight models and production systems. But this reasoning is flawed for a key reason: existing research does not quantify the kinds of information most relevant to a copyright infringement claim [28]. This work typically shows that on average models memorize some portion of their training data; it does not show how much models memorize specific works.
@@
To address this gap, we investigate the extent to which open-weight models memorize verbatim content from copyrighted books. For our experiments, we use the Books3 dataset, a (now notorious) torrented corpus of nearly 200,000200{,}000 books [47, 104, 76]. This dataset is known to be included in the training data of several LLMs [9, 127, 128, 54], and has been a focus of copyright litigation [104, 70]. To gauge whether a particular sequence of text is memorized, we leverage a recent probabilistic extraction method [60] that computes the risk of extracting that sequence at generation time (Section 2).
@@
Extraction risk matters for copyright because the practicability of producing memorized data in a model’s outputs may bear on infringement claims (Section 3). Building on this connection, we develop and validate a procedure that uses probabilistic extraction to surface as much memorization as possible within a book; it uses a sliding window to reveal how much and where memorization occurs, and quantifies the risk that a model will reproduce verbatim sequences in its outputs (Sections 45). We then run thousands of experiments on $50$ books and $17$ LLMs (Section 6). Our results show that, while most models do not memorize most books either in whole or in part, some models memorize some books entirely, such as the first Harry Potter book [106] and 1984 [100]. Following from these results, we show how to deterministically reconstruct a near-perfect copy of the first Harry Potter using Llama 3.1 70B (Section 7). Last, we discuss the significant implications of these findings for copyright cases, though not ones that unambiguously favor either side. Our results complicate defendants’ fair use story, but they also complicate plaintiffs’ efforts to bring class action lawsuits (Section 8). In summary:
@@
  • Section 2. We provide relevant background on memorization and extraction risk [60].
@@ 2 Quantifying memorization through extraction
Extraction and memorization are related concepts, but differ in subtle ways. While extraction refers to reconstructing specific training data in a model’s generated outputs, memorization is broader: it involves reconstructing specific training data by examining the model through any means [29, 28].
Successful extraction of a piece of training data in a model’s generation is evidence of memorization of training data inside the model [28, 15, 114, 76, 45]. This is uncontroversial in machine learning (ML). As Carlini explains, when a sufficiently large and unique piece of training data is extracted, “the only possible explanation is that the model has somewhere internally stored [that piece of training data]” [15]. As Cooper and Grimmelmann note, “in order to be able to extract memorized content from a model at generation time, that memorized content must be encoded in the model’s parameters. There is nowhere else it could be. A model is not a magical portal that pulls fresh information from some parallel universe into our own” [28, p. 25].
@@ 2.1 Probabilistic discoverable extraction
Figure 1: Generating the exact completion of a quote from The Great Gatsby [46] with Llama 1 30B.
@@
Extraction of training data in a model’s outputs—rather than direct inspection of a model’s parameters—is the most common form of evidence used to quantify memorization in LLMs [94, 60, 101, 17, 74, 138]. The standard metric in many frontier model release reports [48, 121, 54] measures verbatim discoverable extraction: given a sequence of text from the training data, split it into a prefix and target suffix, prompt the model with the prefix, and deem extraction to be successful if the LLM reproduces the suffix exactly [16, 18, 93]. On its own, this type of method is well-suited to completion models that continue the prompt text (as in Figure 1), rather than instruction-tuned chatbots that reply in a conversational style. We therefore only study completion models, referring to these models by name, version, and size. For an example, the middle row of Figure 1 shows verbatim extraction of a quote from The Great Gatsby [46] using Llama 1 30B (Appendix A.1).
Discoverable extraction reduces memorization to a binary (yes-or-no) outcome by issuing a single prompt to the model and checking whether the deterministic, greedy-decoded generation matches the target suffix. But LLMs are probabilistic: under realistic, non-deterministic decoding schemes (e.g., top-$k$ sampling), a model can produce different outputs in response to the same prompt, just as in Figure 1. As Cooper and Grimmelmann [28] note, a copyright-relevant claim is therefore more likely to be concerned with how often a target suffix can be extracted, i.e.:
  • a model, when (a) given a particular type of input, will (b) produce a particular type of memorized output, (c) with a particular probability. That probability could be …1% … [or] … 35% …. The issue for copyright law …is what to do …with the fact that element (c)—the probabilistic element—is inescapable [28].
We elaborate on copyright-specific aspects of this point in Sections 3 and 8. For now, we note that Hayes et al. [60] introduce a metric called probabilistic discoverable extraction (probabilistic extraction, for short), which captures this notion. For a training sequence $\bm{z}$} of length a+ja{+}j, one prompts with the $a$-token prefix ${\bm{z}}_{1:a}$ and computes the probability $p_{\bm{z}}\in[0,1]$ of the LLM $\theta$ generating the exact $j$-token suffix ${\bm{z}}_{a+1:a+j}$ when using decoding scheme $\phi$. That is,
@@
In Figure 2, we plot $p_{\bm{z}}$ for the “careless people” quote using top-$k$ (T=1T{=}1, k=40k{=}40) decoding with Llama 1 models, which were trained on The Great Gatsby [127]. The probabilities for this sequence are enormous. In general, $p_{{\bm{z}}}$ is a product of many numbers between 0 and $1$, so even relatively short suffixes should have small values—tending toward 0 as sequence length increases. For scale, a $50$-token suffix with $p_{\bm{z}}\!=\!0.5\%$ means that each token has a very high conditional probability of approximately $90\%$ ($0.5\%\!=\!0.005\!\approx\!0.9^{50}$, i.e., the geometric mean of the token conditional probabilities is $0.005^{\frac{1}{50}}\!\approx\!0.9\!=\!90\%$). [^2] A $100$-token suffix where each token has conditional probability $90\%$ has $p_{\bm{z}}=0.9^{100}\!\approx\!0.0027\%$—about $250\times$ less likely than a $50$-token sequence with the same per-token conditional probabilities. The exception is when the model assigns unusually high probabilities to each token in the sequence. For sequences in the model’s training data, this almost always happens only when the model has memorized those sequences.
@@ 2.2 Interpreting target suffix probabilities as extraction risk
We can therefore interpret $p_{{\bm{z}}}$ as a measure of extraction risk [60]. For a fixed decoding scheme $\phi$ and prefix from the training data, this enables cross-model comparisons: if $p_{{\bm{z}}}^{(A)}>p_{{\bm{z}}}^{(B)}$, then model $A$ will, in expectation, reproduce the target suffix more often than model $B$ (when prompted with that same prefix). For instance, all of the probabilities in Figure 2 exceed $\tau_{\text{min}}\!=\!0.1\%$: they indicate that every Llama 1 model memorized the “careless people” quote in Figure 1. But the quote is not as easily extractable from each model. Llama 3.1 30B exhibits the highest extraction risk for the quote: in expectation, it would take fewer than $3$ prompts (i.e., $\nicefrac{{1}}{{35.2\%}}\approx 2.84$) with the prefix to extract the suffix.
@@
The dozens of pending copyright lawsuits concerning LLMs [21] generally present three interrelated issues: (1) whether training an LLM on copyrighted material is permitted fair use (i.e., limited use of the copyrighted material can, under certain circumstances, be used without permission from the copyright owner); (2) whether the model itself is a copy or derivative work (technical terms in copyright law) of the works on which it is trained (Section 8); and (3) whether the model reproduces copyrighted material in its outputs. Some suits present only one issue, while others present all three questions. Some suits are based on content owned by a single company, while others are class action lawsuits purporting to represent all book authors. [^4] There are also other copyright issues [76, 111, 108, 52, 119, 30], for example, whether novel outputs of generative-AI systems are themselves copyrightable [76, 129]. These lawsuits have primarily focused on (1) training and on (3) model outputs, and the two decisions so far have held that training an LLM can be fair use, albeit with significant limitations [98, 99]. They also raise issues about the acquisition, storage, or use of shadow libraries of copyrighted material as training data, such as Books3 [104, 76].
@@
Our work does not speak to the first issue; training and fair use have been discussed extensively elsewhere [79]. Rather, our findings affect the other two issues: we show that copies are sometimes made in the model itself (i.e., memorized) and those copies can sometimes be generated as outputs (i.e., extracted). Because many LLMs are known to have been trained on Books3, we use books from this dataset to study memorization and extraction of copyrighted material (Sections 46).
@@
Memorization and copies inside the model.  If an LLM memorizes all or a substantial portion of a copyrighted work (near-)verbatim, the LLM itself may be or contain an infringing copy or derivative work [76, 28]. This is because memorized training data are encoded inside the model (Sections 12). Others have argued that encoding the work in the form of model weights satisfies the technical definition of “copy” in the U.S. Copyright Act [28]. In a technical sense, that copy may be of interest to copyright even if there is no practical way to access or extract it. Like the copies of copyrighted works used as training data during training, that copy in model might be fair use, but the analysis of fair use would look somewhat different than internal use of copyrighted works in training processes. This is particularly true for open-weight models like DeepSeek and Llama models, which are not merely used internally by their developers, but are themselves shared externally with others (Section 8).
@@
Extracting copies in model outputs.  Because some memorized works in the training data are extractable, LLMs can sometimes output copies of those works to users [29, 76, 62, 77, 34, 109, 20]. Works with higher extraction probabilities exhibit greater extraction risk; it takes fewer attempts to produce an output that is a verbatim copy of such a work (Section 2.2). That output at generation time will be judged separately from the model itself, and is less likely to be a fair use [62, 76].
@@
Extraction is inherently probabilistic under realistic decoding schemes (Section 2), so the relevant question is not only whether a copy could ever appear for some prompt, but also how often that copy appears in order to be “meaningful”  [28]. When we measure suffix probabilities $p_{\bm{z}}$ that are very close to $0\%$, we do not count those sequences as extracted; one could perhaps understand such generations as instances of “a monkey at the typewriter” [10, 125] (Section 5). Very large $p_{\bm{z}}$ (e.g., $90\%$) indicates near-certain reproduction; this is clearly extraction (and therefore evidence of memorization). However, most cases are not this clear, and it is not immediately obvious where to draw a line in the enormous range between these extremes. To avoid false positives, we use a conservative threshold $\tau_{\min}\!=\!0.1\%$ for scientific evidence of extraction (Sections 2.2 & 5). But for copyright law and policy, the practicability of extraction may require a different bar: if reproducing a work just once would require thousands of prompts, the real-world risk may be relatively limited, even if nonzero (Section 8).

### v4

@@ 1 Introduction
In the dozens of copyright lawsuits over training large language models (LLMs), both sides tend to present opposing, simplified interpretations of the underlying technology. Plaintiffs—such as book authors—say that LLMs are giant copy machines that store (infringing) copies of their specific works and recombine them in outputs [kadrey]. Defendants—typically tech companies—claim that LLMs merely contain non-expressive “statistical correlations” [concordanthropic], and therefore do not copy any particular creative work. We show that the situation is more complicated than either side suggests.
@@
During training, an LLM processes pieces of text in its training data and learns probabilistic relationships across the dataset as a whole. This informal intuition aligns with the “statistical correlations” that defendants describe as encoded in an LLM’s weights. However, sometimes, for some training data, these learned patterns are not so high-level or abstract; rather, they are highly specific [cooper2024files]. For reasons that are not fully understood, certain training data are not so transformed by training: they are “memorized” by the model—encoded in some form inside the LLM’s weights [carlini2023quantifying, lee2023talkin, cooper2024files, carlini2025blog]. Such memorized content can sometimes be “extracted” later; they can be reproduced in an LLM’s outputs at generation time [carlini2021extracting, hayes2025measuringmemorizationlanguagemodels, nasr2023scalable].
@@
Models do not memorize all of their training data, but they do memorize some of it. An LLM may therefore memorize parts of some plaintiffs’ copyrighted works, but not others’. When such memorization occurs, it involves copying that may have significant consequences for ongoing copyright litigation. Extraction generates a “copy” of training data, but it also demonstrates the existence of a “copy” of that training data is memorized inside the model itself. The model being a “copy” (in a technical sense that copyright cares about) has important implications. Notably, a model—not just extracted training data—could be deemed an infringing “copy” of the training data it has memorized [cooper2024files, lee2023talkin]. Copyright law offers the destruction of infringing materials as a remedy. So, just as courts have ordered the destruction of bootleg DVDs, they could order the destruction of infringing models [lee2023talkin, pamscience, townsend2024deletion].
@@
When seeking such remedies, plaintiffs often cite technical papers that demonstrate memorization by extracting training data from open-weight models and production systems. But this reasoning is flawed for a key reason: existing research does not quantify the kinds of information most relevant to a copyright infringement claim [cooper2024files]. This work typically shows that on average models memorize some portion of their training data; it does not show how much models memorize specific works.
@@
To address this gap, we investigate the extent to which open-weight models memorize verbatim content from copyrighted books. For our experiments, we use the Books3 dataset, a (now notorious) torrented corpus of nearly 200,000200{,}000 books [gao2020pile, reisnerbooks3, lee2023talkin]. This dataset is known to be included in the training data of several LLMs [biderman2023pythia, touvron2023llamaopenefficientfoundation, llama2, llama3], and has been a focus of copyright litigation [reisnerbooks3, kadreyamendedconsolidated]. To gauge whether a particular sequence of text is memorized, we leverage a recent probabilistic extraction method [hayes2025measuringmemorizationlanguagemodels] that computes the risk of extracting that sequence at generation time (Section 2).
@@
Extraction risk matters for copyright because the practicability of producing memorized data in a model’s outputs may bear on infringement claims (Section 3). Building on this connection, we develop and validate a procedure that uses probabilistic extraction to surface as much memorization as possible within a book; it uses a sliding window to reveal how much and where memorization occurs, and quantifies the risk that a model will reproduce verbatim sequences in its outputs (Sections 4 & 5). We then run thousands of experiments on $50$ books and $17$ LLMs (Section 6). Our results show that, while most models do not memorize most books either in whole or in part, some models memorize some books entirely, such as the first Harry Potter book [hp1] and 1984 [1984]. Following from these results, we show how to deterministically reconstruct a near-perfect copy of the first Harry Potter using Llama 3.1 70B (Section 7). Last, we discuss the significant implications of these findings for copyright cases, though not ones that unambiguously favor either side. Our results complicate defendants’ fair use story, but they also complicate plaintiffs’ efforts to bring class action lawsuits (Section 8). In summary:
@@
  • Section 2. We provide relevant background on memorization and extraction risk [hayes2025measuringmemorizationlanguagemodels].
@@ 2 Quantifying memorization through extraction
Extraction and memorization are related concepts, but differ in subtle ways. While extraction refers to reconstructing specific training data in a model’s generated outputs, memorization is broader: it involves reconstructing specific training data by examining the model through any means [cooper2023report, cooper2024files].
Successful extraction of a piece of training data in a model’s generation is evidence of memorization of training data inside the model [cooper2024files, carlini2025blog, schwarzschild2024rethinking, lee2023talkin, feldman2020mem]. This is uncontroversial in machine learning (ML). As Carlini explains, when a sufficiently large and unique piece of training data is extracted, “the only possible explanation is that the model has somewhere internally stored [that piece of training data]” [carlini2025blog]. As Cooper and Grimmelmann note, “in order to be able to extract memorized content from a model at generation time, that memorized content must be encoded in the model’s parameters. There is nowhere else it could be. A model is not a magical portal that pulls fresh information from some parallel universe into our own” [cooper2024files, p. 25].
@@ 2.1 Probabilistic discoverable extraction
Figure 1: Generating the exact completion of a quote from The Great Gatsby [gatsby] with Llama 1 30B.
@@
Extraction of training data in a model’s outputs—rather than direct inspection of a model’s parameters—is the most common form of evidence used to quantify memorization in LLMs [nasr2025scalable, hayes2025measuringmemorizationlanguagemodels, prashanth2024recite, carlini2023extracting, lee2022dedup, zhang2023counterfactual]. The standard metric in many frontier model release reports [reid2024gemini, gemma2, llama3] measures verbatim discoverable extraction: given a sequence of text from the training data, split it into a prefix and target suffix, prompt the model with the prefix, and deem extraction to be successful if the LLM reproduces the suffix exactly [carlini2021extracting, carlini2023quantifying, nasr2023scalable]. On its own, this type of method is well-suited to completion models that continue the prompt text (as in Figure 1), rather than instruction-tuned chatbots that reply in a conversational style. We therefore only study completion models, referring to these models by name, version, and size. For an example, the middle row of Figure 1 shows verbatim extraction of a quote from The Great Gatsby [gatsby] using Llama 1 30B (Appendix A.1).
Discoverable extraction reduces memorization to a binary (yes-or-no) outcome by issuing a single prompt to the model and checking whether the deterministic, greedy-decoded generation matches the target suffix. But LLMs are probabilistic: under realistic, non-deterministic decoding schemes (e.g., top-$k$ sampling), a model can produce different outputs in response to the same prompt, just as in Figure 1. As cooper2024files note, a copyright-relevant claim is therefore more likely to be concerned with how often a target suffix can be extracted, i.e.:
  • a model, when (a) given a particular type of input, will (b) produce a particular type of memorized output, (c) with a particular probability. That probability could be …1% … [or] … 35% …. The issue for copyright law …is what to do …with the fact that element (c)—the probabilistic element—is inescapable [cooper2024files].
We elaborate on copyright-specific aspects of this point in Sections 3 and 8. For now, we note that hayes2025measuringmemorizationlanguagemodels introduce a metric called probabilistic discoverable extraction (probabilistic extraction, for short), which captures this notion. For a training sequence $\bm{z}$} of length a+ja{+}j, one prompts with the $a$-token prefix ${\bm{z}}_{1:a}$ and computes the probability $p_{\bm{z}}\in[0,1]$ of the LLM $\theta$ generating the exact $j$-token suffix ${\bm{z}}_{a+1:a+j}$ when using decoding scheme $\phi$. That is,
@@
In Figure 2, we plot $p_{\bm{z}}$ for the “careless people” quote using top-$k$ (T=1T{=}1, k=40k{=}40) decoding with Llama 1 models, which were trained on The Great Gatsby [touvron2023llamaopenefficientfoundation]. The probabilities for this sequence are enormous. In general, $p_{{\bm{z}}}$ is a product of many numbers between 0 and $1$, so even relatively short suffixes should have small values—tending toward 0 as sequence length increases. For scale, a $50$-token suffix with $p_{\bm{z}}\!=\!0.5\%$ means that each token has a very high conditional probability of approximately $90\%$ ($0.5\%\!=\!0.005\!\approx\!0.9^{50}$, i.e., the geometric mean of the token conditional probabilities is $0.005^{\frac{1}{50}}\!\approx\!0.9\!=\!90\%$). [^2] A $100$-token suffix where each token has conditional probability $90\%$ has $p_{\bm{z}}=0.9^{100}\!\approx\!0.0027\%$—about $250\times$ less likely than a $50$-token sequence with the same per-token conditional probabilities. The exception is when the model assigns unusually high probabilities to each token in the sequence. For sequences in the model’s training data, this almost always happens only when the model has memorized those sequences.
@@ 2.2 Interpreting target suffix probabilities as extraction risk
We can therefore interpret $p_{{\bm{z}}}$ as a measure of extraction risk [hayes2025measuringmemorizationlanguagemodels]. For a fixed decoding scheme $\phi$ and prefix from the training data, this enables cross-model comparisons: if $p_{{\bm{z}}}^{(A)}>p_{{\bm{z}}}^{(B)}$, then model $A$ will, in expectation, reproduce the target suffix more often than model $B$ (when prompted with that same prefix). For instance, all of the probabilities in Figure 2 exceed $\tau_{\text{min}}\!=\!0.1\%$: they indicate that every Llama 1 model memorized the “careless people” quote in Figure 1. But the quote is not as easily extractable from each model. Llama 3.1 30B exhibits the highest extraction risk for the quote: in expectation, it would take fewer than $3$ prompts (i.e., $\nicefrac{{1}}{{35.2\%}}\approx 2.84$) with the prefix to extract the suffix.
@@
The dozens of pending copyright lawsuits concerning LLMs [chatgptiseating] generally present three interrelated issues: (1) whether training an LLM on copyrighted material is permitted fair use (i.e., limited use of the copyrighted material can, under certain circumstances, be used without permission from the copyright owner); (2) whether the model itself is a copy or derivative work (technical terms in copyright law) of the works on which it is trained (Section 8); and (3) whether the model reproduces copyrighted material in its outputs. Some suits present only one issue, while others present all three questions. Some suits are based on content owned by a single company, while others are class action lawsuits purporting to represent all book authors. [^4] There are also other copyright issues [lee2023talkin, pamscience, sag2023safety, goodyear2025infringement, sobelstyle, cooper2024unlearning], for example, whether novel outputs of generative-AI systems are themselves copyrightable [lee2023talkin, uscopyrightofficereport2]. These lawsuits have primarily focused on (1) training and on (3) model outputs, and the two decisions so far have held that training an LLM can be fair use, albeit with significant limitations [kadreyjudgment, bartzjudgment]. They also raise issues about the acquisition, storage, or use of shadow libraries of copyrighted material as training data, such as Books3 [reisnerbooks3, lee2023talkin].
@@
Our work does not speak to the first issue; training and fair use have been discussed extensively elsewhere [lemley2021fairlearning]. Rather, our findings affect the other two issues: we show that copies are sometimes made in the model itself (i.e., memorized) and those copies can sometimes be generated as outputs (i.e., extracted). Because many LLMs are known to have been trained on Books3, we use books from this dataset to study memorization and extraction of copyrighted material (Sections 46).
@@
Memorization and copies inside the model.  If an LLM memorizes all or a substantial portion of a copyrighted work (near-)verbatim, the LLM itself may be or contain an infringing copy or derivative work [lee2023talkin, cooper2024files]. This is because memorized training data are encoded inside the model (Sections 1 & 2). Others have argued that encoding the work in the form of model weights satisfies the technical definition of “copy” in the U.S. Copyright Act [cooper2024files]. In a technical sense, that copy may be of interest to copyright even if there is no practical way to access or extract it. Like the copies of copyrighted works used as training data during training, that copy in model might be fair use, but the analysis of fair use would look somewhat different than internal use of copyrighted works in training processes. This is particularly true for open-weight models like DeepSeek and Llama models, which are not merely used internally by their developers, but are themselves shared externally with others (Section 8).
@@
Extracting copies in model outputs.  Because some memorized works in the training data are extractable, LLMs can sometimes output copies of those works to users [cooper2023report, lee2023talkin, henderson2023fair, lee2024talkinshort, cyphert2024memlaw, sag2024fairness, charlesworth2025]. Works with higher extraction probabilities exhibit greater extraction risk; it takes fewer attempts to produce an output that is a verbatim copy of such a work (Section 2.2). That output at generation time will be judged separately from the model itself, and is less likely to be a fair use [henderson2023fair, lee2023talkin].
@@
Extraction is inherently probabilistic under realistic decoding schemes (Section 2), so the relevant question is not only whether a copy could ever appear for some prompt, but also how often that copy appears in order to be “meaningful” [cooper2024files]. When we measure suffix probabilities $p_{\bm{z}}$ that are very close to $0\%$, we do not count those sequences as extracted; one could perhaps understand such generations as instances of “a monkey at the typewriter” [borelmonkey, simpsons] (Section 5). Very large $p_{\bm{z}}$ (e.g., $90\%$) indicates near-certain reproduction; this is clearly extraction (and therefore evidence of memorization). However, most cases are not this clear, and it is not immediately obvious where to draw a line in the enormous range between these extremes. To avoid false positives, we use a conservative threshold $\tau_{\min}\!=\!0.1\%$ for scientific evidence of extraction (Sections 2.2 & 5). But for copyright law and policy, the practicability of extraction may require a different bar: if reproducing a work just once would require thousands of prompts, the real-world risk may be relatively limited, even if nonzero (Section 8).